For events A1, A2, A2 (not necessarily disjoint) define disjoint events B1, B2, B3 (i.e. Bi ∩ Bj = \varnothingfor i 6= j) such that, A1 ∪ A2 ∪ A3 = B1 ∪ B2 ∪ B3.
Answer #1

A, UA, UNO A, uf(An- a.) – [(4,143) – (Aim di = (A2-417 - 4412 443 74A10 Azn hafa ar (A2-A) - (ARNABAU (ANAL NAB) = B, and Az